# Learning large softmax mixtures with warm start EM

### v1

@@ B.3.3 Proof of Lemma 4: concentration inequality of the EM-updates within the specified neighborhood
and taking the union bounds over $k\in[K]$. ∎
B.3.4 Proofs of Lemmas 7, 5 and 6
@@
This bound, the inequality
[…]
To prove (91), we have
${|\pi_{{\bm{\omega}}}(X_{j})-\pi_{{\bm{\omega}}^{\prime}}(X_{j})|\over\pi_{{\bm{\omega}}}(X_{j})}$ $\leq{1\over\pi_{{\bm{\omega}}}(X_{j})}\sum_{k=1}^{K}\left(|\alpha_{k}-\alpha_{k^{\prime}}|A_{{\bm{\theta}}_{k}}(X_{j})+\alpha_{k}^{\prime}\left|A_{{\bm{\theta}}_{k}}(X_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})\right|\right)$
The proof is complete. ∎
Proof of Lemma 6.
${1\over\sum_{\ell=1}^{p}e^{X_{\ell}^{\top}\bar{\bm{\theta}}^{*}}}\sup_{v\in\mathbb{S}^{L-1}}\sum_{j=1}^{p}e^{X_{j}^{\top}\bar{\bm{\theta}}^{*}}v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})^{\top}Ev$
$=\sup_{v\in\mathbb{S}^{L-1}}\left\{{v^{\top}\mathrm{I}_{\bar{\bm{\theta}}^{*}}v\over N_{\bar{\bm{\theta}}^{*}}}+(v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})^{2}-{2v^{\top}\mathrm{II}_{\bar{\bm{\theta}}^{*}}\over N_{\bar{\bm{\theta}}^{*}}}(v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})\right\}.$
By repeating the arguments in the proof of Lemma 13 to control the deviation $\mathrm{I}_{\bar{\bm{\theta}}^{*}}/N_{\bar{\bm{\theta}}^{*}}-\mathbb{E}[\mathrm{I}_{\bar{\bm{\theta}}^{*}}]/\mathbb{E}[N_{\bar{\bm{\theta}}^{*}}]$ and $\mathrm{II}_{\bar{\bm{\theta}}^{*}}/N_{\bar{\bm{\theta}}^{*}}-\mathbb{E}[\mathrm{II}_{\bar{\bm{\theta}}^{*}}]/\mathbb{E}[N_{\bar{\bm{\theta}}^{*}}]$ via Lemmas 8, 9 and 10, together with the expressions in Lemma 11, one can deduce that
$\sup_{v\in\mathbb{S}^{L-1}}\left\{{v^{\top}\mathrm{I}_{\bar{\bm{\theta}}^{*}}v\over N_{\bar{\bm{\theta}}^{*}}}+(v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})^{2}-{2v^{\top}\mathrm{II}_{\bar{\bm{\theta}}^{*}}\over N_{\bar{\bm{\theta}}^{*}}}(v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})\right\}$
$\leq\sup_{v\in\mathbb{S}^{L-1}}\left\{v^{\top}\Sigma v+(v^{\top}\Sigma\bar{\bm{\theta}}^{*})^{2}-2(v^{\top}\Sigma\bar{\bm{\theta}}^{*})(v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})+(v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})^{2}\right\}$
$\qquad+\sup_{v\in\mathbb{S}^{L-1}}|v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}}|~{}C\overline{\sigma}\sqrt{L+\log(p)\over p^{1-\delta}}+C\overline{\sigma}^{2}\sqrt{L+\log(p)\over p^{1-\delta}}$ (119)
with probability at least $1-4p^{1-s}$. Furthermore, since
$v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}}-v^{\top}\Sigma{\bm{\theta}}_{k}^{*}$ $=v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}}^{*}+v^{\top}\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}}-A_{{\bm{\theta}}_{k}}^{*})-{v^{\top}\mathbb{E}[\mathrm{II}_{{\bm{\theta}}_{k}^{*}}]\over\mathbb{E}[N_{{\bm{\theta}}_{k}^{*}}]}$ by Lemma 11
[…]
By repeating arguments of proving Lemma 13, invoking Lemma 9 and Lemma 10 and applying the bound in (B.3.1) to control $v^{\top}\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}^{*}}(X_{j})-A_{{\bm{\theta}}_{k}}(X_{j}))$, with probability at least $1-4p^{1-s}$,
$\left|v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}}-v^{\top}\Sigma{\bm{\theta}}_{k}^{*}\right|$ $\lesssim\overline{\sigma}\sqrt{L+\log(p)\over p^{1-\delta}}+\delta_{0}\|\mathbf{X}\|_{\infty,2}\sqrt{\|\mathbf{X}^{\top}\mathrm{diag}(A({\bm{\theta}}_{k}^{*}))\mathbf{X}\|_{\mathrm{op}}}$
$\lesssim\overline{\sigma}\sqrt{L+\log(p)\over p^{1-\delta}}+\delta_{0}\|\mathbf{X}\|_{\infty,2}\sqrt{\overline{\sigma}^{2}+\overline{\sigma}^{4}\|{\bm{\theta}}_{k}^{*}\|_{2}^{2}}$ by Lemma 11
holds uniformly over $v\in\mathbb{S}^{L-1}$. Note the above display also implies
$\sup_{v\in\mathbb{S}^{L-1}}|v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}}|$ $\lesssim\overline{\sigma}^{2}\|{\bm{\theta}}_{k}^{*}\|_{2}^{2}+\overline{\sigma}\sqrt{L+\log(p)\over p^{1-\delta}}+\delta_{0}\|\mathbf{X}\|_{\infty,2}\sqrt{\overline{\sigma}^{2}+\overline{\sigma}^{4}\|{\bm{\theta}}_{k}^{*}\|_{2}^{2}}$
[…]
In conjunction with (B.3.4), we conclude that
$\sup_{v\in\mathbb{S}^{L-1}}\left\{{v^{\top}\mathrm{I}_{\bar{\bm{\theta}}^{*}}v\over N_{\bar{\bm{\theta}}^{*}}}+(v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})^{2}-{2v^{\top}\mathrm{II}_{\bar{\bm{\theta}}^{*}}\over N_{\bar{\bm{\theta}}^{*}}}(v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})\right\}$
$\lesssim\sup_{v\in\mathbb{S}^{L-1}}\left\{v^{\top}\Sigma v+(v^{\top}\Sigma\bar{\bm{\theta}}^{*})^{2}-2(v^{\top}\Sigma\bar{\bm{\theta}}^{*})(v^{\top}\Sigma{\bm{\theta}}_{k}^{*})+(v^{\top}\Sigma{\bm{\theta}}_{k}^{*})^{2}\right\}$ (120)
$\qquad+C(\overline{\sigma}^{2}B^{2}+\overline{\sigma})\overline{\sigma}\sqrt{L+\log(p)\over p^{1-\delta}}+C\overline{\sigma}^{2}B^{2}~{}\delta_{0}\|\mathbf{X}\|_{\infty,2}\sqrt{\overline{\sigma}^{2}\left(1+\overline{\sigma}^{2}B^{2}\right)}$
with probability at least $1-4p^{1-s}$. Since the terms in line (B.3.4) equal to
$\sup_{v\in\mathbb{S}^{L-1}}\left\{v^{\top}\Sigma v+\left(v^{\top}\Sigma\bar{\bm{\theta}}^{*}-v^{\top}\Sigma{\bm{\theta}}_{k}^{*}\right)^{2}\right\}\leq\overline{\sigma}^{2}+{\overline{\sigma}^{4}\over 4}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2},$
using $L+\log p\leq cp^{1-\delta}$ completes the proof for any fixed ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$.
Finally, inspecting the proof reveals that only bounding the term $v^{\top}\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}}-A_{{\bm{\theta}}_{k}}^{*})$ depends on ${\bm{\omega}}$. Since its bound in (B.3.1) holds uniformly over ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$, the claimed result for fixed $a\neq k$ follows. The proof is completed by the union bounds argument over $a,k\in[K]$. ∎
Proof of Lemma 7.
[…]
Further recall from Lemma 11 that for any fixed ${\bm{\theta}}\in\mathbb{R}^{L}$,
$\mathbb{E}[N_{{\bm{\theta}}}]=p\exp({\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2).$
We work on the event
$\mathcal{E}_{p}=\bigcap_{a,k=1}^{K}\left\{{1\over 2}\leq{N_{{\bm{\theta}}_{k}^{*}}\over\mathbb{E}[N_{{\bm{\theta}}_{k}^{*}}]}\leq 2,~{}{1\over 2}\leq{N_{\bar{\bm{\theta}}_{ak}^{*}}\over\mathbb{E}[N_{\bar{\bm{\theta}}_{ak}^{*}}]}\leq 2\right\}.$ (121)
By part (a) of Lemma 12, after taking the union bound over $a,k\in[K]$, we find that this event has probability $1-4K^{2}p^{1-s}$. Since $A_{{\bm{\theta}}_{k}^{*}}(X_{j})=e^{X_{j}^{\top}{\bm{\theta}}_{k}^{*}}/N_{{\bm{\theta}}_{k}^{*}}$, we further write
${\lesssim}{(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(X_{j}^{\top}({\bm{\theta}}_{a}^{*}+{\bm{\theta}}_{k}^{*})/2\right)/p\over\exp\left({\bm{\theta}}_{k}^{*\top}\Sigma{\bm{\theta}}_{k}^{*}/2+X_{j}^{\top}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})/2\right)+\exp\left({\bm{\theta}}_{a}^{*\top}\Sigma{\bm{\theta}}_{a}^{*}/2-X_{j}^{\top}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})/2\right)}$
on the event $\mathcal{E}_{p}$. From the identity
${1\over 2}{\bm{\theta}}_{a}^{*\top}\Sigma{\bm{\theta}}_{a}^{*}+{1\over 2}{\bm{\theta}}_{k}^{*\top}\Sigma{\bm{\theta}}_{k}^{*}={1\over 4}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})^{\top}\Sigma({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})+\bar{\bm{\theta}}_{ak}^{*\top}\Sigma\bar{\bm{\theta}}_{ak}^{*},$
the denominator of the inequality above is
$\exp\left({1\over 2}{\bm{\theta}}_{k}^{*\top}\Sigma{\bm{\theta}}_{k}^{*}+X_{j}^{\top}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})/2\right)+\exp\left({1\over 2}{\bm{\theta}}_{a}^{*\top}\Sigma{\bm{\theta}}_{a}^{*}-{1\over 2}X_{j}^{\top}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})\right)$
$=\exp\left({1\over 4}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})^{\top}\Sigma({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})+\bar{\bm{\theta}}_{ak}^{*\top}\Sigma\bar{\bm{\theta}}_{ak}^{*}-{1\over 2}{\bm{\theta}}_{a}^{*\top}\Sigma{\bm{\theta}}_{a}^{*}+{1\over 2}X_{j}^{\top}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})\right)$
$\qquad+\exp\left({1\over 2}{\bm{\theta}}_{a}^{*\top}\Sigma{\bm{\theta}}_{a}^{*}-{1\over 2}X_{j}^{\top}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})\right)$
$\geq 2\exp\left({1\over 8}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})^{\top}\Sigma({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})+{1\over 2}\bar{\bm{\theta}}_{ak}^{*\top}\Sigma\bar{\bm{\theta}}_{ak}^{*}\right).$
It then follows that
${\alpha_{a}^{*}\alpha_{k}^{*}\over A_{\bar{\bm{\theta}}_{ak}^{*}}(X_{j})}{A_{{\bm{\theta}}_{a}^{*}}(X_{j})A_{{\bm{\theta}}_{k,u}}(X_{j})\over\pi_{{\bm{\omega}}}(X_{j})}$
$\lesssim(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(-{1\over 8}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})^{\top}\Sigma({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})\right){\exp(X_{j}^{\top}\bar{\bm{\theta}}_{ak}^{*})\over p\exp(\bar{\bm{\theta}}_{ak}^{*\top}\Sigma\bar{\bm{\theta}}_{ak}^{*}/2)}{1\over A_{\bar{\bm{\theta}}_{ak}^{*}}(X_{j})}$
$\lesssim(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(-{\underline{\sigma}^{2}\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right)$
on the event $\mathcal{E}_{p}$, which completes the proof. ∎
@@
Appendix C Concentration inequalities for quantities related with the random embedding matrix

### v2

@@ C.6 Proof of Theorem 4
Invoking Lemma 21, Lemma 16 and Lemma 14 and taking union bounds over $k\in[K]$ give that
$\sup_{v\in\mathbb{S}^{L-1}}|\bar{m}_{r}(v)-m_{r}(v)|\lesssim r^{r/2}\sqrt{L\log(p)\over p}+[r(L+\log p)]^{r/2}{L\log(p)\over p^{1-\delta/2}}$ (123)
with probability at least $1-p^{-s}$.
Since the same argument applies to prove the bounds for the errors of the mixed-moments, we omit the proof. ∎
C.7 Extension to $\mu=\mathcal{N}_{L}(0,\Sigma)$
When $X\sim\mathcal{N}_{L}(0,\Sigma)$, one can still use
$h_{r}(X)$ $=H_{r}(X^{\top}v),$ (124)
$h_{r1;i}(X)$ $=H_{r}(X^{\top}v)(X^{\top}w_{i})\,.$ (125)
Let $U:=\Sigma^{-1/2}X\sim\mu_{0}=\mathcal{N}_{L}(0,\mathbf{I}_{L})$. Then, for any generic ${\bm{\theta}}\in\mathbb{R}^{L}$, and given $v\in\mathbb{R}^{L}$
$\mathbb{E}_{\mu}\left[H_{r}(X^{\top}v)\exp(X^{\top}{{\bm{\theta}}})\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{{\bm{\theta}}})\right]$ $=$ ${\mathbb{E}_{\mu_{0}}\left[H_{r}(U^{\top}u)\exp(U^{\top}{\bar{{\bm{\theta}}}})\right]\over\mathbb{E}_{\mu_{0}}\left[\exp(U^{\top}{\bar{{\bm{\theta}}}})\right]}\ =\ (u^{\top}\bar{{{\bm{\theta}}}})^{r}\ =\ (v^{\top}\Sigma{{\bm{\theta}}})^{r}$ (126)
with $u:=\Sigma^{1/2}v$ and $\bar{{{\bm{\theta}}}}:=\Sigma^{1/2}{{\bm{\theta}}}$, where the second equality holds by Lemma 11, by the construction of $h_{r}$, since $U$ is a standard Gaussian on $\mathbb{R}^{L}$. Thus, if the procedure of Section 3.3 is applied relative to functions given by (124), but $X\sim\mathcal{N}_{L}(0,\Sigma)$, then $\bar{{\bm{\theta}}}_{k}$ approximates $\Sigma{{\bm{\theta}}}_{k}^{*}$, for each $k\in[K]$. One immediately has
$\|\Sigma^{-1}\bar{{\bm{\theta}}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\leq{1\over\underline{\sigma}^{2}}\|\bar{{\bm{\theta}}}_{k}-\Sigma{\bm{\theta}}_{k}^{*}\|_{2}.$
C.8 Proof of Theorem 5
Proof.
From Proposition 2, it suffices to show
$\max_{r<2K}|\widehat{m}_{r}-\bar{m}_{r}|+\max_{r<K,2\leq i\leq L}|\widehat{m}_{r1;i}-\bar{m}_{r1:i}|=\mathcal{O}_{\mathbb{P}}(\sqrt{\log(L)/N}).$
For every fixed integer $r<2K$, by an application of Chebyshev’s inequality, for any $t>0$,
$\mathbb{P}\left(|\widehat{m}_{r}-\bar{m}_{r}|\geq\frac{t}{\sqrt{N}}\mid\mathbf{X}\right)$$\leq$ $\frac{\mathbb{E}[h_{r}^{2}(Y)\mid\mathbf{X}]}{t^{2}}.$
Since the quantity
$\mathbb{E}\left[h_{r}^{2}(Y)\mid\mathbf{X}\right]$ $=$ $\sum_{k=1}^{K}\alpha_{k}^{*}{{1\over p}\sum_{j=1}^{p}h_{r}^{2}(X_{j})\exp(X_{j}^{\top}{\bm{\theta}}_{k}^{*})\over{1\over p}\sum_{i=1}^{p}\exp(X_{i}^{\top}{\bm{\theta}}_{k}^{*})}$
has $\mu$ a.s. limit, by taking the union bounds over $r<2K$ with $K=\mathcal{O}(1)$, we conclude that for large $p$,
$\max_{r<2K}|\widehat{m}_{r}-\bar{m}_{r}|=\mathcal{O}_{\mathbb{P}}(\sqrt{\log(K)/N}).$
Similar arguments can be used to bound $\max_{r<K,2\leq i\leq L}|\widehat{m}_{r1;i}-\bar{m}_{r1:i}|$. ∎
Remark 11.
[…]
In particular, the proof of Proposition 2 reveals that its conclusion holds when each atom of $\widehat{\rho}_{1}$ and $\rho_{1}^{*}$ is at least $\Delta_{1}/2$ away from all but $\ell^{\prime}=1$ other atom (itself). However, if $\ell^{\prime}>1$, Proposition 4 in Wu and Yang (2020) shows that we cannot expect a parametric rate in the estimation of ${\bm{\theta}}_{k}$, even in one dimension, as display (109) in the proof then becomes
$W_{1}(\widehat{\rho}_{1},\rho_{1}^{*})\leq 2K\left({2K4^{2K-1}2^{2K-\ell^{\prime}-1}\over\Delta_{1}^{2K-\ell^{\prime}-1}}\right)^{1\over\ell^{\prime}}\|\widehat{\bm{m}}-{\bm{m}}\|_{2}^{1\over\ell^{\prime}},$ (127)
a rate that will be inherited by $|\theta_{1k}^{*}-\widehat{\theta}_{1k}|$, for each $k$, via (105). In the worst case, when $\ell^{\prime}=2K-1$, we obtain $W_{1}(\widehat{\rho}_{1},\rho_{1})\lesssim K\|\widehat{\bm{m}}-\bm{m}\|_{2}^{1/(2K-1)},$ by Proposition 1 in Wu and Yang (2020). Thus, although consistent estimation of the softmax mixture parameters will continue to hold when the atoms are distinct, but not well separated, neither the estimation of ${\bm{\theta}}_{k}^{*}$ nor that of ${\bm{\alpha}}^{*}$ can be expected to follow a parametric decay rate. This is confirmed by our simulation results in Section 4.
C.9 Proof of Proposition 3
Proof.
By definition, it suffices to bound
$\left\|\widehat{\Gamma}-\bar{\Gamma}\right\|_{\mathrm{op}}=\left\|{1\over N}\sum_{\ell=1}^{N}W_{\ell}\right\|_{\mathrm{op}},$
where we write
$W_{\ell}:={\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}-\mathbb{E}\left[{\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}\mid\mathbf{X}\right]\in\mathbb{R}^{L\times L}.$
To invoke the matrix-valued Bernstein’s inequality in Lemma 31, note that, by using (55),
$\max_{\ell\in[N]}\left\|W_{\ell}\right\|_{\mathrm{op}}\leq 2\max_{\ell\in[N]}\left\|{\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}\right\|_{\mathrm{op}}\leq 2C\|\mathbf{X}\|_{\infty,2}^{2}.$ (128)
while
$\left\|\sum_{\ell=1}^{N}\mathbb{E}[W_{\ell}^{2}]\right\|_{\mathrm{op}}$ $\leq N\left\|\mathbb{E}\left[{\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}{\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}\mid\mathbf{X}\right]\right\|_{\mathrm{op}}$
$\leq N\left\|\mathbb{E}\left[{\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}\mid\mathbf{X}\right]\right\|_{\mathrm{op}}C\|\mathbf{X}\|_{\infty,2}^{2}$ by (128)
$\leq N\left\|\bar{\Gamma}\right\|_{\mathrm{op}}C\|\mathbf{X}\|_{\infty,2}^{2}$ by (53)
$\leq N\left(\|\bar{\Gamma}-\Gamma\|_{\mathrm{op}}+\|\Gamma\|_{\mathrm{op}}\right)C\|\mathbf{X}\|_{\infty,2}^{2}$
[…]
Invoking Lemma 31 with $\sigma^{2}=C^{\prime}N\|\mathbf{X}\|_{\infty,2}^{2}$, $U=2C\|\mathbf{X}\|_{\infty,2}^{2}$ and $t=C^{\prime\prime}\sqrt{N\log(N)}\|\mathbf{X}\|_{\infty,2}$ yields that, on the event $\mathcal{E}_{\Gamma}(\epsilon_{p}^{\prime})$,
$\|\widehat{\Gamma}-\bar{\Gamma}\|_{\mathrm{op}}\lesssim\|\mathbf{X}\|_{\infty,2}\sqrt{\log N\over N},$
with probability at least $1-14\exp\left(-C^{\prime\prime}\log(N)+\log(L)\right).$ The proof is complete. ∎
C.10 Proof of 3
Proof.
For the case $\Sigma=\mathbf{I}_{L}$, we have
$\|\bar{\Gamma}-\Gamma\|_{\mathrm{op}}=\sup_{v\in\mathbb{S}^{L-1}}|\bar{m}_{r}(v)-m_{r}(v)|$
so that invoking part (2) of Theorem 6 with $r=2$ gives that for any $\delta>0$, $s\geq 1$ and $p\geq p_{0}(B,\delta,s)$, $\mathbb{P}\{\mathcal{E}_{\Gamma}(\epsilon_{p}^{\prime})\}\geq 1-p^{-s}$ with
$\epsilon_{p}^{\prime}\lesssim\sqrt{L\log(p)\over p}+{(L+\log p)L\log(p)\over p^{1-\delta/2}}.$
The claim thus follows by recalling $\mathcal{E}_{2}$ in (120).
For the general case $\mathcal{N}_{L}(0,\Sigma)$, it is easy to see that $\widehat{\Gamma}$ and $\Gamma$ are rescaled version of their counterparts for $\Sigma=\mathbf{I}_{L}$ (written as $\widehat{\Gamma}_{0}$ and $\Gamma_{0}$) in the sense that
$\widehat{\Gamma}-\Gamma=\Sigma^{-1/2}(\widehat{\Gamma}_{0}-\Gamma_{0})\Sigma^{-1/2}.$
The claim thus follows immediately. ∎
C.11 Proof of Lemma 2
Proof.
Pick any $k\neq k^{\prime}$. We first bound from below
$|v^{\top}{\bm{\theta}}_{k}^{*}-v^{\top}{\bm{\theta}}_{k^{\prime}}^{*}|={|({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})^{\top}\widehat{V}\widehat{V}^{\top}u|\over\|\widehat{V}\widehat{V}^{\top}u\|_{2}}={|({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})^{\top}\widehat{V}(\widehat{V}^{\top}u)|\over\|\widehat{V}^{\top}u\|_{2}}.$
$\mathbb{P}\left\{|v^{\top}{\bm{\theta}}_{k}^{*}-v^{\top}{\bm{\theta}}_{k^{\prime}}^{*}|<\|\widehat{V}({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})\|_{2}~t\right\}<t\sqrt{K}.$
To bound from above $\|\widehat{V}^{\top}({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})\|_{2}$, recall that $\widehat{V}\in\mathbb{O}_{L\times K}$ denotes the left leading eigenvectors of $\widehat{\Gamma}$. It then follows that
$\|\widehat{V}^{\top}({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})\|_{2}^{2}$ $=({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})^{\top}\widehat{V}\widehat{V}^{\top}({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})$
$=\|{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*}\|_{2}^{2}-({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})^{\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})$
$\geq\Delta^{2}-2{\bm{\theta}}_{k}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top}){\bm{\theta}}_{k}^{*}-2{\bm{\theta}}_{k^{\prime}}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top}){\bm{\theta}}_{k^{\prime}}^{*}.$
Notice that
${\bm{\theta}}_{k}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top}){\bm{\theta}}_{k}^{*}$ $=\sup_{u\in\mathbb{S}^{L-1}}u^{\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top}){\bm{\theta}}_{k}^{*}{\bm{\theta}}_{k}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})u$
$\leq{1\over\underline{\alpha}}\sup_{u\in\mathbb{S}^{L-1}}u^{\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})\sum_{a=1}^{K}\alpha_{a}^{*}{\bm{\theta}}_{a}^{*}{\bm{\theta}}_{a}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})u$
Therefore, since the event $C^{\prime}\|\mathbf{X}\|_{\infty,2}\sqrt{\log N/N}+\epsilon_{p}^{\prime}\leq\underline{\alpha}\Delta^{2}$ and Proposition 3 imply
$\Delta^{2}\underline{\alpha}\geq 8\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}},$ (129)
we obtain that for all $t\geq 0$,
$\mathbb{P}\left\{|v^{\top}{\bm{\theta}}_{k}^{*}-v^{\top}{\bm{\theta}}_{k^{\prime}}^{*}|<{t\Delta\over 2\sqrt{K}}\right\}<t.$
By taking the union bounds over $k,k^{\prime}\in[K]$ with $k\neq k^{\prime}$, the proof is complete. ∎
C.12 Proof of Lemma 3
We first give a proof of Eq. 61 for completeness.
Proof of Eq. 61.
Fix ${\bm{\theta}}^{*}\in\mathbb{S}^{L-1}$. For arbitrary $\delta\in[0,1)$, Tkocz (2012) gives that for any $v$ uniformly drawn from $\mathbb{S}^{L-1}$,
$\mathbb{P}\left\{v^{\top}{\bm{\theta}}^{*}\geq\delta\right\}\leq e^{-L\delta^{2}/2}.$
It thus implies that, for any $\varepsilon\in(0,1)$,
$\mathbb{P}\left\{v_{i}^{\top}{\bm{\theta}}^{*}<\delta,~\text{for all }i\in[m]\right\}$ $~\leq~\left(1-e^{-L\delta^{2}/2}\right)^{m}$
$~\leq~e^{-me^{-L\delta^{2}/2}}$ $\text{by }1-x\leq e^{-x}$
$~\leq~\varepsilon$
provided that
$m\geq\exp(L\delta^{2}/2)\log(1/\varepsilon).$
Recall that for any $v,{\bm{\theta}}^{*}\in\mathbb{S}^{L-1}$
$\|v-{\bm{\theta}}^{*}\|_{2}\leq\delta_{0}\quad\iff\quad v^{\top}{\bm{\theta}}^{*}\geq 1-{\delta_{0}^{2}\over 2}.$
The proof is complete by taking $\delta=1-\delta_{0}^{2}/2$ above and using $(1-\delta_{0}^{2}/2)^{2}\leq 2-\delta_{0}^{2}$. ∎
@@ C.12 Proof of Lemma 3
Recall that ${\bm{\theta}}_{k}^{*}\in\mathbb{S}^{L-1}$. For any $i\in[m]$ with $v_{i}$ defined in (60), one has
$\|v_{i}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}$ $=2-2v_{i}^{\top}{\bm{\theta}}_{k}^{*}$
$=2-{2u_{i}^{\top}\widehat{V}\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\over\|\widehat{V}^{\top}u_{i}\|_{2}}$
$=2-{2u_{i}^{\top}\widehat{V}\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\over\|\widehat{V}^{\top}u_{i}\|_{2}\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2}}+{2u_{i}^{\top}\widehat{V}\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\over\|\widehat{V}^{\top}u_{i}\|_{2}\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2}}(1-\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2})$
$\leq\left\|{\widehat{V}^{\top}u_{i}\over\|\widehat{V}^{\top}u_{i}\|_{2}}-{\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\over\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2}}\right\|_{2}^{2}+2\left(1-\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2}\right).$
Using the arguments in the proof of Proposition 3, one has
$1-\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2}^{2}={\bm{\theta}}_{k}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top}){\bm{\theta}}_{k}^{*}\leq{2\over\underline{\alpha}}\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}:=\delta_{\Gamma}.$
The proof follows by applying (61) to the first term with $\delta^{2}=\delta_{0}^{2}-2\delta_{\Gamma}$. ∎
@@
Appendix D Concentration inequalities for quantities related with the random embedding matrix